OpenStudy (anonymous):

Help please. Will give medal. How do i convert 12.71 (g) of NcHCO3 to NaCl ? I need help getting the answer.

OpenStudy (ipwnbunnies):

You need to have the chemical equation first.

OpenStudy (anonymous):

37.06 (g) - 24.35 (g) = 12.71 (g) of NcHCO3

OpenStudy (cwrw238):

NaHCO3 + HCl = nacl + H2O + CO2

OpenStudy (cwrw238):

* NaCl

OpenStudy (anonymous):

ok

OpenStudy (cwrw238):

plug in the relative atomic masses of the elements to find the proportions in which they react

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Can someone please help me figure out what the correct answer is

OpenStudy (phi):

can you find the atomic mass of one molecule of \[ Na H C O_3\]?